OpenStudy (anonymous):

State the coordinates of the y-intercept of the graph of -4x - 6y = 13. Your answer should be an ordered pair.

OpenStudy (marihelenh):

You need to get it into slope-intercept form which is y=mx+b. You would: 1) Add 4x to both sides. -6y=4x+13 2) Divide each side by -6 to get the equation into slope-intercept form. y=-2/3x-13/6 The y intercept is the b in y=mx+b. So, your y intercept is -13/6. Wherever the y intercept is, that means that x is equal to zero. So, you ordered pair is (x,y)=(0,-13/6) \[(0,\frac{ -13 }{ 6 })\]

OpenStudy (usukidoll):

to find the y-intercept, we have to set our x to 0 and then just solve for y... and that's already been done and it's the correct answer. Well done @marihelenh!